Rick always likes to consume a cup of tea with a serving of noodles at any given time of the day. He earns a monthly allowance of 150. A cup of tea cost him Rs5 and serving of noodles cost Rs 10. Given this what will be Rick’s optimal consumption bundle?


1
Expert's answer
2021-10-08T09:57:16-0400

"I = 150\\

P_t = 5\\

P_n =10"

Here, t is denoted as Tea and n is denoted as noodles.

Therefore,

"I=(P_t)t+(P_n)n\\150=5t+10n ........... (i)\\Now,\\Slope \space of\space IC = −MRS_{t,n}=−\frac{MU_t}{MU_n}\\ MU_t=n\\ MU_n=t\\Slope\space of\space IC=−\frac{t}{n}\\Slope\space of\space Budget\space line=−\frac{Pt}{Pn}\\ =−\frac{5}{10}\\ =−\frac{1}{2}"

Slope of IC = Slope of Budget Line

"−\frac{t}{n}=−\frac{1}{2}\\\frac{t}{n}=\frac{1}{2}\\2t=n ........... (ii)"


Now by putting the value of n in equation (i) we get,

"150=5t+10(2t)\\

150=5t+20t\\

25t=150\\

t=6\\

Or \space t^*=6"

Now, putting the value of t in equation (ii) we get,

"2(6)=n\\

n=12\\

Or \space n^*=12"


Hence Rick’s optimal consumption bundle is t* = 6 and n* = 12